Real interest rate (percent per year) Supply of loanable funds (2005 dollars) Demand for loanable funds (2005 dollars) 5 2,000 5,000 7 3,000 4,000 9 4,000 3,000 11 5,000 2,000 a) Draw the demand and supply curves. b) What is the equilibrium real interest rate? c) What is equilibrium investment? Equilibrium saving? d) Describe the situation in Dream Island's loanable funds market when the real interest rate is 10 percent. Is there a shortage of loanable funds? A surplus of loanable funds? e) Describe the situation in Dream Island's capital market when the real interest rate is 6 percent. Is there a shortage of loanable funds? A surplus of loanable funds? 


1
Expert's answer
2021-08-05T14:15:54-0400

a)

At equilibrium, demand = supply

"-500 + 500I = 7,500 - 500I\\

I = 8\\

F = 3,500"



b) 8%


c) 3,500


d) At a interest rate of 10%, there is demand of loanable funds of"F = 7,500 - 500 \times 10 = 2,500"

Supply of funds at "I = 10\space is\space -500 + 500 \times 10 = 4,500"

surplus of loanable funds "= 4,500 - 2,500 = 2,000" .


e) At a interest rate of 6%, there is demand of loanable funds of "F = 7,500 - 500 \times6 = 4,500"

Supply of funds at "I = 6 is -500 + 500 \times 6 = 2,500"

There is shortage of loanable funds of "4,500 - 2,500 = 2,000"